OpenStudy (shadowcain):

A blueprint for a house shows the height of the house to be 3 1/2 inches. The actual house is 21 1/3 feet tall. What is the unit rate in feet per inch? A. 7/128 foot per inch B. 6 2/21 feet per inch C. 21/128 foot per inch D. 2 2/7 feet per inch

jimthompson5910 (jim_thompson5910):

they want you to divide `21 1/3` by `3 1/2`

jimthompson5910 (jim_thompson5910):

first you'll need to convert the mixed numbers to improper fractions example: 3 1/2 = 3+1/2 3 1/2 = 3*(2/2)+1/2 3 1/2 = 6/2+1/2 3 1/2 = 7/2

OpenStudy (shadowcain):

Don't you have to convert feet into inches first?

jimthompson5910 (jim_thompson5910):

no, just divide the two numbers without converting notice how the rate is "feet per inch", so you'll have "feet" over "inches"

OpenStudy (shadowcain):

Thanks

jimthompson5910 (jim_thompson5910):

you're welcome